Diffusion Morning Report 01/23/04

ted berg tberg at snf.stanford.edu
Fri Jan 23 07:06:13 PST 2004


Good Morning All,
    Looking pretty good this morning.  Had a few issues with SiGe 
yesterday hopefully they are resolved.  Epi pump was low on oil and had
kicked off , it is back up now- doing leak check.   TEOS2 is moving 
along.  Colby is Taming the THERMCOs.  Still waiting for them to be put 
on Coral
Everything else looking OK.

Diffusion 	Date: 	1/23/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	TEOS Conversion 	 
Tylan 8-SiGe 	yellow 	leak n2/ar-reset  zero 	1.2/30
Tylan 9-Poly 	Up 	  	24.8/30
Tylan 10-Nitride 	Up 	  	25.9/30
Tylan 11-BPSG-LTO 	Up 	  	16.45/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	17.5/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	Checked this morning 	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 225- B(c4) 
on line


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040123/ce0ee013/attachment.html>


More information about the diffusion mailing list